# PPJudge: Towards Human-Aligned Assessment of Artistic Painting Process

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; 2025; Dublin, Irelandccs: Computing methodologies Computer vision tasksccs: Applied computing Fine arts
@@ 1. Introduction
  • We propose a novel Painting Process Assessment Model called PPAM, which incorporates temporal-aware position encoding and a heterogeneous MoE architecture.
@@ 2.3. Sequence Image Representation
Sequence image representation has evolved significantly over the past decades, covering both image and video domains. Convolutional neural networks (CNNs) were long the dominant architecture for visual representation learning (He et al., 2016; Huang et al., 2017). With the rise of Transformer-based models (Dosovitskiy et al., 2021), vision transformers (ViTs) have become popular image encoders. DeiT (Touvron et al., 2021) introduces knowledge distillation to enable data-efficient ViT training, while Swin Transformer (Liu et al., 2021a) proposes hierarchical attention with local windows for improved efficiency and scalability. To further enhance image encoders, researchers explore sparse and lightweight architectures. Switch Transformers (Fedus et al., 2022) introduces sparse routing to reduce computation, which is widely adopted in large-scale models (Team, 2024a; DeepSeek-AI, 2024).
Compared to image encoders, video encoders additionally capture temporal dynamics across frames. Early approaches extended 2D CNNs to 3D to model spatial and temporal features (Carreira and Zisserman, 2017). SlowFast Networks (Feichtenhofer et al., 2019) introduces dual pathways to capture both slow and fast temporal patterns. Recently, many works have adapted Transformers to video understanding. ViViT (Arnab et al., 2021) proposes factorized attention strategies for efficient spatio-temporal modeling, while TimeSformer (Bertasius et al., 2021) decouples spatial and temporal attention to reduce complexity. Video Swin Transformer (Liu et al., 2022) extends hierarchical attention to the temporal dimension. Self-supervised video pretraining has also gained attention. VideoMAE (Tong et al., 2022; Wang et al., 2023) leverages masked video modeling to learn strong temporal-spatial representations from unlabeled videos.
Image encoders process sequential images by simply aggregating the embeddings of individual frames, which fails to effectively capture inter-frame relationships. In contrast, video encoders model temporal dynamics but require re-encoding the entire frame sequence at each step, making real-time assessment impractical. In this paper, we propose an efficient design based on an image encoder augmented with key-value caching. Our method supports incremental updates, avoiding redundant computation and significantly improving inference efficiency.
@@ 3.1. Painting Collection
Following (Jiang et al., 2024) and (Liu et al., 2024), we recruit hundreds of art and drawing majors as participants aged between 22 and 30. Before painting, each participant is given either a random prompt or a reference painting. Participants can create freely based on the prompt or replicate the reference as faithfully as possible. The entire painting process is recorded as a video focused solely on the canvas. Since participants often pause to observe or think, the videos contain many redundant frames with little visual change. Moreover, painting durations vary significantly across individuals. To address these issues, we first downsample the video by extracting one frame every $k$ frames. We then compute the visual difference between consecutive frames based on pixel-wise L2 distance. A frame is selected as a key frame if its visual difference from the last selected key frame exceeds a threshold $\tau$. To adapt to varying painting speeds and session durations, we dynamically adjust $\tau$ based on the total visual change across the video. Finally, we enforce minimum and maximum limits on the number of extracted key frames to ensure consistent representation quality.
@@ 3.3. Painting Annotation
Painting Consistency (Consis.) evaluates whether the final painting accurately reflects the core themes and elements of the given prompt or reference image, maintaining strong semantic and visual similarity. A higher score indicates stronger alignment between the final painting and the reference.
@@
Painting Depth assesses whether the painting process exhibits progressive refinement through the gradual introduction of new techniques or complex visual elements (e.g., color schemes, composition, and light-shadow treatments). This aspect is evaluated in terms of detail, color, and composition. Detail depth (D. D.) measures the incremental enrichment of visual details and their contribution to expressive power; Color depth (Col. D.) evaluates whether the use of color becomes more sophisticated over time, enhancing visual impact; Composition depth (Com. D.) assesses whether compositional structures grow in complexity and contribute to the overall visual richness.
In total, the painting process is assessed using eight attributes on a scale from 1 to 10. Specifically, each painting’s drawing process is evaluated by multiple experts, and each attribute receives multiple scores. The final score for each attribute is computed as the average of these scores. This approach helps mitigate the impact of subjectivity in the evaluation process.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3754960ccs: Computing methodologies Computer vision tasksccs: Applied computing Fine arts
@@ 1. Introduction
  • We propose a novel painting process assessment model called PPJudge, which incorporates temporal-aware position encoding and a heterogeneous MoE architecture.
@@ 2.3. Sequence Image Representation
Sequence image representation has evolved significantly over the past decades, covering both image and video domains. Convolutional neural networks (CNNs) were long the dominant architecture for visual representation learning (He et al., 2016; Huang et al., 2017). Recently, vision transformers (ViTs) have become popular image encoders. DeiT (Touvron et al., 2021) introduces knowledge distillation to enable data-efficient ViT training, while Swin Transformer (Liu et al., 2021a) proposes hierarchical attention with local windows for improved efficiency and scalability. To further enhance image encoders, researchers explore sparse and lightweight architectures (Fedus et al., 2022).
Compared to image encoders, video encoders additionally capture temporal dynamics across frames. Early approaches extended 2D CNNs to 3D to model spatial and temporal features (Carreira and Zisserman, 2017). Recently, many works have adapted Transformers to video representation. ViViT (Arnab et al., 2021) proposes factorized attention strategies for efficient spatio-temporal modeling, while TimeSformer (Bertasius et al., 2021) decouples spatial and temporal attention to reduce complexity. Video Swin Transformer (Liu et al., 2022) extends hierarchical attention to the temporal dimension. Self-supervised video pretraining has also gained attention. VideoMAE (Tong et al., 2022; Wang et al., 2023) utilizes masked video modeling to learn strong temporal-spatial representations from unlabeled videos.
Image encoders process sequential images by simply aggregating the embeddings of individual frames, which fails to effectively capture inter-frame relationships. In contrast, video encoders model temporal dynamics but require re-encoding the entire frame sequence at each step. In this paper, we propose an efficient design based on an image encoder augmented with key-value caching. Our method supports incremental updates, avoiding redundant computation and significantly improving inference efficiency.
@@ 3.1. Painting Collection
Following (Jiang et al., 2024) and (Liu et al., 2024), we recruit hundreds of art and drawing majors as participants aged between 22 and 30. Each participant is given either a random prompt or a reference painting. Participants can create freely based on the prompt or replicate the reference faithfully. The entire painting process is recorded as a canvas-focused video. Since participants often pause to observe or think, videos contain redundant frames, and painting durations vary widely. To address this, we sample one frame every $k$ frames and compute visual differences between consecutive frames. A frame is selected as a key frame if its visual difference from the last selected key frame exceeds a threshold $\tau$, which is dynamically adjusted based on the total visual change across the video. To ensure consistent representation, we further constrain the number of key frames within predefined bounds.
@@ 3.3. Painting Annotation
Painting Consistency (Consis.) evaluates whether the final painting accurately reflects the core themes and elements of the given prompt or reference image, maintaining strong semantic and visual similarity.
@@
Painting Depth assesses whether the painting process exhibits progressive refinement through the gradual introduction of new techniques or complex visual elements. This aspect is evaluated in terms of detail, color, and composition. Detail depth (D. D.) measures the incremental enrichment of visual details and their contribution to expressive power; Color depth (Col. D.) evaluates whether the use of color becomes more sophisticated over time, enhancing visual impact; Composition depth (Com. D.) assesses whether compositional structures grow in complexity and contribute to the overall visual richness.
In total, the painting process is assessed using eight attributes on a scale from 1 to 10. Specifically, each painting’s drawing process is evaluated by multiple experts, and each attribute receives multiple scores. The final score for each attribute is computed as the average of these scores.
@@ 6. Conclusion
Acknowledgements.
The authors wish to acknowledge the support from NSSFC under Grant 22ZD05, NNSFC under Grant 62472178, and the Natural Science Foundation of Shanghai Municipality, China under Grant 24ZR1418300.